OpenStudy (anonymous):

factorise the following expression if possible by completing the square: 3x^2 -6x-5

OpenStudy (ahsome):

First, you need to take 3 as common, as the "a" value needs to be 1 Rewrite Equation: \[3(x^2-2x+\frac{-5}{-4})\] Simplify: \[3(x^2-2x+\frac{5}{4})\] To compelete the square, you need to first do this: \[(\frac{b}{2})^2\] Sub in the b value \[(\frac{1}{2})^2\] This gives: \[0.25\] Then put this value between the "b" and "c" value, making sure to take away \[3(x^2-2x+0.25-0.25+\frac{5}{4})\] You can group the first 3 values \[3((x-0.25)^2-0.25+\frac{5}{4})\] Simplify \[3((x-0.25)^2+1)\] Expand: \[3(x-0.25)^2+3\] Square root, then square the 3 \[3(x-0.25)^2+(\sqrt{3})^2\] Then use the Difference of Square Law \[3(x-0.25-\sqrt{3})(x-0.25+\sqrt{3})\] Done, phew ;)

OpenStudy (ahsome):

@TL9800 Hope this helped. Since this is such a winded question, need someone like @iambatman to clarrify

OpenStudy (anonymous):

mhm

OpenStudy (ahsome):

Whoops, replace the 4 by 3. Not sure where that came from ;)

OpenStudy (anonymous):

\[3x^2-6x-5\] you can group the terms containing x, leaving a blank spot to add constant in later \[(3x^2-6x+~~~)-5 \] I think you forgot to do this part @Ahsome factor of a 3 now \[3(x^2-2x+~~~~)-5\] \[3(\frac{ -2 }{ 2 })^2 =3\] \[3(x^2-2x+1)-5-3 \implies 3(x^2-2x+1)-8 \implies 3(x-1)^2-8\]

OpenStudy (ahsome):

That's another way. This was the way I learnt in school. @TL9800 Remember to change 4 to 3, error there ;)

OpenStudy (anonymous):

thank uoi

OpenStudy (ahsome):

No Problem mate ;)